A baseball player hits a ball such that as it leaves the bat, it has speed of 40 meters per second and has an angle of 440. What is the time that the ball reaches the highest point of its flight?


1
Expert's answer
2021-08-23T04:49:14-0400

"w= \Delta \theta \Delta t"

"\Delta \theta" is the angular rotation

w is the velocity

"\Delta t = time"

"40 m/s = 440^o Xt\\
t = 0.091 sec"